Graduate Research Assistant, NPAC, Syracuse University, Syracuse, N.Y. 13244 email : zeynep@npac.syr.edu

May 26, 1999

Prof. Kishan G. Mehrotra, Program Director CIS, Syracuse University Syracuse, N.Y. 13244

PhD Progress Report - Spring 1999:

Grade  Course Name                                   Instructor
--------------------------------------------------------------------------
A      CIS623 Structured Programming and Semantics   H. Blair
A-     CIS720 Introductory Category Theory           S. Older
A      CSE791 Distributed Object Technologies        J. Fawcett
Co-Author
---------
1- "WebHLA as Integration Platform for FMS and other Metacomputing
   Application Domains", In Proceedings of the DoD HPC Users Group Conference,
   Monterey, CA, June 8-15, 1999
2- "Using WebHLA to Integrate HPC FMS Modules with Web/Commodity based
   Distributed Object Technologies of CORBA, Java, COM and XML,
   in Proceedings of the Advanced Simulation Technologies Conference (ASTC99),
   San Diego, CA, Apr 11-15, 1999.

Sincerely,

Zeynep Odcikin-Ozdemir